# Finite Elements with weighted bases for the fractional Laplacian

### v1

@@ 1 Introduction and main results
Finite element methods for problem (1.1) have been studied using piece-wise linear basis functions on polygonal domains (see, for instance, Acosta & Borthagaray [4, 5]). The reported convergence rates for these methods, in $s$-fractional Sobolev-Slobodeckij norm, are at best approximately $h^{\frac{1}{2}}$, where $h$ is the diameter of the mesh triangles. The sublinear convergence rates are due to the regularity of the solution $u$, which is at most $u\in H^{s+\frac{1}{2}-\varepsilon}(\Omega)$ for small $\varepsilon>0$ (we refer the reader to Proposition 3.6 and Proposition 3.11 in [5]). This reduced regularity arises from the fact that, even in smooth domains, the solution $u$ behaves like $d_{\Omega}^{s}$ in a neighbourhood of the boundary $\partial\Omega$, where $d_{\Omega}(x)=\mathrm{dist}(x,\mathbb{R}^{d}\setminus\Omega)$. We refer to [2] for a recent result of this type and Section 2.7 in [30] for a summary of the related regularity results. The low regularity of $u$ makes higher convergence rates difficult to obtain using a standard piece-wise linear basis that is normally used for solutions that are at least weakly differentiable once. Higher convergence rates of order $h^{\frac{3}{2}-s}$ in $s$-fractional Sobolev-Slobodeckij norm using piece-wise linear basis functions have been proved in [9] only under the assumption of higher regularity for the solution like for example $u(x)=(1-|x|^{2})_{+}$ with $\Omega=B_{1}$, which imposes to work with a sign changing right-hand side $f$ (see more comments on this matter in Section 3).
@@
Our approach builds on the observation that, intuitively, if $\delta$ denotes a smooth approximation of the distance function $d_{\Omega}$, then the quotient $u/\delta^{s}$ exhibits significantly better regularity up to the boundary, provided that both $f$ and the domain $\Omega$ are sufficiently smooth. This regularity theory has been developed through a series of contributions [48, 35, 46, 47, 45, 2, 3] (see [30] for a good survey on the matter). Motivated by this fact, we construct a tailored finite element basis that captures the improved regularity of $u/\delta^{s}$, thereby achieving higher-order convergence rates than those available in the existing literature. The idea, that will be detailed in (1.12), is to use a basis of the form $\delta^{s}\times($piece-wise linear functions)). One of the key advantages of the finite element method is its effectiveness in approximating smooth functions and our method takes advantage of this well-known fact in order to approximate $u/\delta^{s}$ and subsequently, knowing $\delta^{s}$, the solution $u$ itself. This motivates our strategy of approximating $u/\delta^{s}$ rather than $u$ directly. The advantage of this approach is that we can obtain improved convergence rates, without requiring higher regularity of $u$, but rather $f$ and $\Omega$ sufficiently smooth, which is often the case in applications. Assuming enough regularity on the right-hand side $f$ and the domain $\Omega$, we obtain convergence rates of order $h^{2-s}$ in $s$-fractional Sobolev-Slobodeckij norm, improving all the previously reported results in the literature. Using Aubin-Nitsche duality, we obtain $L^{2}$-rates of higher than our $H^{s}$-rates.
At the end of the paper we provide numerical experiments that precisely match the $H^{s}$-rate of order $h^{2-s}$, and that suggest the $L^{2}$-rate is of order $h^{2}$.
@@ 1.1 Preliminaries and functional setting
$H^{s}(\mathbb{R}^{d})\coloneq\left\{w\in L^{2}(\mathbb{R}^{d}):[w]_{H^{s}(\mathbb{R}^{d})}<\infty\right\},\qquad H^{s}_{0}(\Omega)\coloneq\overline{C_{c}^{\infty}(\Omega)}^{H^{s}},$
endowed with the norm $\|w\|_{H^{s}(\mathbb{R}^{d})}\coloneq\|w\|_{L^{2}(\mathbb{R}^{d})}+[w]_{H^{s}(\mathbb{R}^{d})}$. Due to the fractional Poincaré inequality
$\|u\|_{L^{2}(\Omega)}\leq C[u]_{H^{s}({\mathbb{R}^{d}})},\qquad\forall u\in H^{s}_{0}(\Omega),$ (1.4)
the $\cdot]_{H^{s}}$ seminorm is actually an equivalent norm of $H^{s}_{0}(\Omega)$ (for further details we refer the reader to [23, 12]). When $\partial\Omega\in C^{0,1}$ we have the alternative characterisation (see, e.g., [23])
$H^{s}_{0}(\Omega)=\{w\in H^{s}({\mathbb{R}^{d}})\text{ such that }w=0\text{ in }{\mathbb{R}^{d}}\setminus\Omega\}.$ (1.5)
Finally, for $\alpha>0$ we denote by $W^{\alpha,\infty}(\Omega)$ the Hölder space of order $\alpha$
@@
Setting $V\coloneq H_{0}^{s}(\Omega)$, the problem (1.1) can be written in weak form as follows:
@@
Let us also define $N_{h}\coloneq\dim V_{h}$. For the remaining elements, we drop the dependence on $h$ for the sake of the readability. We define $\{x_{i}\}_{i=1}^{N_{h}}$ be the vertexes of all simplexes $K\in\mathcal{T}_{h}$ (including those lying in $\partial\Omega_{h}$). The canonical basis of $\mathrm{PL}(\mathcal{T}_{h})$ is given by the piece-wise linear function $\varphi_{i}$ such that $\varphi_{i}(x_{j})=\delta_{ij}$. The canonical basis of $V_{h}$ is $\phi_{i}=\delta^{s}\varphi_{i}$. When we use this basis, we will refer to the weighted Finite-Elements method, or WFEM for short.
@@
Let $s\in(0,1)$, $\partial\Omega\in C^{1}$ and $\delta$ satisfy ($\textup{A}_{\delta}^{\sigma}$) with $\sigma=1$. Then $V_{h}\subset H^{s}_{0}(\Omega)$.
@@
Due to Remark 1.2, if $\Omega=B_{1}$ and $f=1$, and we pick the distance $\delta(x)=1-|x|^{2}$ we have precisely that $u/\delta^{s}$ is constant. This means that $u\in V_{h}$. Since $V_{h}\subset V$, then $u$ satisfies (1.13), and by uniqueness $u=u_{h}$. The numerical solution is exactly the true solution.
@@ 1.2 Main results
Let $s\in(0,1)$, $\mu\in(s,2]$, $\partial\Omega\in C^{1}$ if $\mu\in(s,1]$, $\partial\Omega\in C^{1,\mu-1}$ if $\mu\in(1,2]$, and $\delta$ satisfy ($\textup{A}_{\delta}^{\sigma}$) with $\sigma=\max\{1,\mu\}$. Let also $u$ be the corresponding solution of (1.6), assume that $u/\delta^{s}\in W^{\mu,\infty}({\Omega})$, and $V_{h}$ be given by (1.12) and $u_{h}$ be the solution of (1.13). Then, we have that
@@
where $C$ does not depend on $h$. In particular, the maximal order $h^{2-s}\lvert\log h\rvert^{\frac{1}{2}}$ is achieved if $\mu=2$, i.e., $\partial\Omega\in C^{1,1}$ and $u/\delta^{s}\in W^{2,\infty}(\Omega)$.
@@
Let $s\in(0,1)$, $\gamma\in(s,3)$ such that $\gamma,\gamma\pm s\not\in\mathbb{N}$, $f\in W^{\gamma-s,\infty}(\Omega)$, $\partial\Omega\in C^{\gamma+1}$, and ($\textup{A}_{\delta}^{\sigma}$) holds with $\sigma=\gamma+1$. Let $u$ be the corresponding solution of (1.6), $V_{h}$ be given by (1.12), and $u_{h}$ be the solution of (1.13). Then, we have that
@@
Notice that in the results above, the constants $C$ depend on all the parameters in the statement $s,\mu,\mu$ (or $\gamma$), the geometric parameters $\Omega,d$, the constants in ($\textup{A}_{\delta}^{\sigma}$), and the uniform constants in the definition of the triangulation $\mathcal{T}_{h}$.
@@ 1.3 Comments on related literature
Our results indicate that the use of weighted basis can substantially improve the convergence rate, achieving up to $h^{2-s}\lvert\log h\rvert^{1/2}$, albeit at the cost of requiring higher regularity of both $f$ and $\Omega$. Furthermore, Theorem 4.9 in [5] demonstrates that when graded meshes are employed, the convergence rate in (1.16) can be improved up to an almost linear rate, proportional to $N^{-\frac{1}{d}}|\log N^{-\frac{1}{d}}|^{1/2}$, for $s\in(1/2,1)$ where in $N$ is the degrees of freedom (recall that in the case of uniform meshes $N_{h}\sim h^{-d}$). In the present work, we piece-wise linearly approximate $u/\delta^{s}$ which sufficiently smooth everywhere in $\Omega$ so graded meshes do not seem to provide a major improvement. This also simplifies the implementation.
@@
Finally, we emphasize that both our work and the previously cited contributions address problem (1.1) involving the fractional Laplacian in its integral form, as defined in (1.2) over the entire space $\mathbb{R}^{d}$. However, the fractional Laplacian admits several distinct definitions (see, e.g., [39, 40]). Suitable localizations of these definitions to domains yield non-equivalent operators. Among these alternative formulations, the spectral fractional Laplacian has received particular attention due to its applications to problem (1.1) and to other equations involving fractional diffusion operators. The solutions of the corresponding Dirichlet problem behave like $u\sim d_{\Omega}$ ([16]), so the main difficulty of our framework is not present. This operator also admits a Caffarelli-Silvestre type extension ([51]) Finite element methods for the extension problem were first introduced in the pioneering work [42], achieving a convergence rate of $\lvert\log N\rvert^{s}N^{-1/(d+1)}$ in the $s$ fractional Sobolev norm corresponding to the problem, where $N$ stands for the degrees of freedom. To deal with computational challenges arising from the extra dimension, see [6]. Subsequently, in [7], the authors obtain exponential $H^{s}$ convergence rates using $hp$-FEM. For other approaches see [10, 21, 20, 52, 17].
@@ 2 Proofs of the main results
For any $\mu\in(0,2]$ let $\Omega$ be such that $\partial\Omega\in C^{0,1}$ if $\mu\in(0,1]$ and $C^{1,\mu-1}$ if $\mu\in(1,2]$, and consider a linear extension operator $E_{\mu}$ satisfying
@@
The existence of such an extension operator can be found in [50, Chapter VI, Theorem 3] and [33, Lemma 6.37]. We point out that in the construction of Stein $E_{\mu}=E_{1}$ for all $\mu\in(0,1]$. Furthermore, $E_{1}:C(\overline{\Omega})\to C_{b}({\mathbb{R}^{d}})$. Lastly, let us introduce the restriction $r_{\Omega}$ such that $r_{\Omega}u=u|_{\Omega}$. We first make the observation that for $C^{1}$ domains the regularized distance function $\delta$ belongs to $H^{s}_{0}(\Omega)$.
@@
Assume that $\partial\Omega\in C^{1}$ and that $\delta\in W^{1,\infty}(\Omega)$ satisfies $c\delta\leq d_{\Omega}\leq C\delta$. Then, $\delta^{s}\in H^{s}_{0}(\Omega)$ for all $s\in(0,1)$.
@@
Since $\Omega\in C^{1}$ we can use the tubular neighbourhood theorem to prove that
$\int_{\Omega}d_{\Omega}(x)^{s-1}\mathrm{d}x<\infty.$
Since $\delta=0$ outside $\Omega$ we can use the characterisation (1.5) of $H_{0}^{s}(\Omega)$ to conclude the result. ∎
@@ 2.1 Proof of convergence rates in $H^{s}$
As usual in the finite elements setting, we rely on the following Cea’s lemma to prove the desired convergence result.
@@
Let $s\in(0,1)$, $\partial\Omega\in C^{1}$ and $\delta$ satisfy ($\textup{A}_{\delta}^{\sigma}$) with $\sigma=1$. Let also $u$ be the corresponding solution of (1.6), $V_{h}$ be given by (1.12), and $u_{h}$ be the solution of (1.13). Then,
@@
We recall that $V_{h}\subset H^{s}_{0}(\Omega)$ as shown in Lemma 1.5. The $\geq$ inequality in (2.2) holds because $u_{h}\in V_{h}$. Let $v_{h}\in V_{h}$. For any $w_{h}\in V_{h}$ we can use the weak formulation to deduce $a(u-u_{h},w_{h})=a(u,w_{h})-a(u_{h},w_{h})=\int_{\Omega}fw_{h}-\int_{\Omega}fw_{h}=0$. Therefore, we have that
@@
In order to build a suitable competitor to apply Cea’s Lemma and obtain the desired orders of convergence, we need to define a suitable interpolant. More precisely, let $I_{h}:C(\mathbb{R}^{d})\to PL(\mathcal{T}_{h})\subset W^{1,\infty}({\Omega}_{h})$ be the piece-wise linear interpolation obtained by sampling on the vertexes of the triangles in $\mathcal{T}_{h}$. We define, for $\mu\in(s,2]$ the map $J_{h}:\delta^{s}W^{\mu,\infty}({\mathbb{R}^{d}})\to V_{h}$ by
@@
Applying the $K$-interpolation results in [41, Theorem 1.6, Example 1.8, result after the proof Example 1.8] that $T_{h}:W^{\alpha,\infty}({\mathbb{R}^{d}})\to W^{\alpha,\infty}({\mathbb{R}^{d}}),$ and we have
@@
First, we apply [19, Theorem 3.1.5] to show the result for some integer $\alpha,\mu$, i.e.,
@@
Let $s\in(0,1)$, $\mu\in(s,2]$, $\partial\Omega\in C^{1}$ if $\mu\in(s,1]$, $\partial\Omega\in C^{1,\mu-1}$ if $\mu\in(1,2]$, and $\delta$ satisfy ($\textup{A}_{\delta}^{\sigma}$) with $\sigma=\max\{1,\mu\}$. Let also $v\in\delta^{s}C(\overline{\Omega})$ such that $v/\delta^{s}\in W^{\mu,\infty}(\overline{\Omega})$. Then, for all $\alpha\in(s,\min\{\mu,1\}]$, we have that
$v-J_{h}v]_{H^{s}(\mathbb{R}^{d})}\leq\frac{C_{1}}{\sqrt{\alpha-s}}h^{\mu-\alpha}\|v/\delta^{s}\|_{W^{\mu,\infty}({\Omega})}.$ (2.7)
@@
$v-J_{h}v]_{H^{s}(\mathbb{R}^{d})}\leq C_{2}h^{\mu-s}\lvert\log h\rvert^{\frac{1}{2}}\|v/\delta^{s}\|_{W^{\mu,\infty}({\Omega})}$ (2.8)
@@
$v-J_{h}v=\delta^{s}E_{\alpha}r_{\Omega}E_{\mu}\frac{v}{\delta^{s}}-\delta^{s}E_{\alpha}r_{\Omega}I_{h}E_{\mu}\frac{v}{\delta^{s}}=\delta^{s}E_{\alpha}r_{\Omega}\left(E_{\mu}\frac{v}{\delta^{s}}-I_{h}E_{\mu}\frac{v}{\delta^{s}}\right)$

### v2

@@ 1 Introduction and main results
Finite element methods for problem (1.1) have been studied using piece-wise linear basis functions on polygonal domains (see, for instance, Acosta & Borthagaray [4, 5]). The reported convergence rates for these methods, in $s$-fractional Sobolev-Slobodeckij norm, are at best approximately $h^{\frac{1}{2}}$, where $h$ is the diameter of the mesh triangles. The sublinear convergence rates are due to the regularity of the solution $u$, which is at most $u\in H^{s+\frac{1}{2}-\varepsilon}(\Omega)$ for small $\varepsilon>0$ (we refer the reader to Proposition 3.6 and Proposition 3.11 in [5]). This reduced regularity arises from the fact that, even in smooth domains, the solution $u$ behaves like $d_{\Omega}^{s}$ in a neighbourhood of the boundary $\partial\Omega$, where $d_{\Omega}(x)=\mathrm{dist}(x,\mathbb{R}^{d}\setminus\Omega)$. We refer to [2] for a recent result of this type and Section 2.7 in [33] for a summary of the related regularity results. The low regularity of $u$ makes higher convergence rates difficult to obtain using a standard piece-wise linear basis that is normally used for solutions that are at least weakly differentiable once. In [9], the authors point out that when $u$ is sufficiently regular and linear up to the boundary (e.g., $u(x)=(1-|x|^{2})_{+}$ with $\Omega=B_{1}$), the higher-order convergence rate $h^{\frac{3}{2}-s}$ in the $s$-fractional Sobolev–Slobodeckij norm can be achieved using piecewise linear basis functions. It is worth noting that such functions $u$ correspond to solutions of problems with a sign-changing right-hand side $f$ (see further discussion in Section 3).
@@
Our approach builds on the observation that, intuitively, if $\delta$ denotes a smooth approximation of the distance function $d_{\Omega}$, then the quotient $u/\delta^{s}$ exhibits significantly better regularity up to the boundary, provided that both $f$ and the domain $\Omega$ are sufficiently smooth. This regularity theory has been developed through a series of contributions [52, 39, 50, 51, 49, 2, 3] (see [33] for a good survey on the matter). Motivated by this fact, we construct a tailored finite element basis that captures the improved regularity of $u/\delta^{s}$, thereby achieving higher-order convergence rates than those available in the existing literature. The idea, that will be detailed in (1.11), is to use a basis of the form $\delta^{s}\times($piece-wise linear functions)). One of the key advantages of the finite element method is its effectiveness in approximating smooth functions and our method takes advantage of this well-known fact in order to approximate $u/\delta^{s}$ and subsequently, knowing $\delta^{s}$, the solution $u$ itself. This motivates our strategy of approximating $u/\delta^{s}$ rather than $u$ directly. The advantage of this approach is that we can obtain improved convergence rates, without requiring higher regularity of $u$, but rather $f$ and $\Omega$ sufficiently smooth, which is often the case in applications. Assuming enough regularity on the right-hand side $f$ and the domain $\Omega$, we obtain convergence rates of order $h^{2-s}$ in $s$-fractional Sobolev-Slobodeckij norm, improving all the previously reported results in the literature. Using Aubin-Nitsche duality, we obtain $L^{2}$-rates of higher than our $H^{s}$-rates. At the end of the paper we provide numerical experiments that precisely match the $H^{s}$-rate of order $h^{2-s}$, and that suggest the $L^{2}$-rate is of order $h^{2}$.
@@ 1.1 Preliminaries and functional setting
$H^{s}(\mathbb{R}^{d})\coloneq\left\{w\in L^{2}(\mathbb{R}^{d}):[w]_{H^{s}(\mathbb{R}^{d})}<\infty\right\},$
endowed with the norm $\|w\|_{H^{s}(\mathbb{R}^{d})}\coloneq\|w\|_{L^{2}(\mathbb{R}^{d})}+[w]_{H^{s}(\mathbb{R}^{d})}$. We also introduce the corresponding energy space
$V\coloneq\{w\in H^{s}({\mathbb{R}^{d}})\text{ such that }w=0\text{ in }{\mathbb{R}^{d}}\setminus\Omega\},$
which is sometimes denoted by $\widetilde{H}^{s}(\Omega)$ in the literature.
Remark 1.1.
It is standard to define the trace spaces for $s\in(0,1)$ as $H^{s}_{0}(\Omega)\coloneq\overline{C_{c}^{\infty}(\Omega)}^{H^{s}}$. When $\partial\Omega\in C^{0,1}$ and $s\neq\tfrac{1}{2}$, we have the alternative characterization $V=H^{s}_{0}(\Omega)$ (see, e.g., [26]). In the case $s=\tfrac{1}{2}$, the space $V$ coincides with the so-called Lions–Magenes space $H^{1/2}_{00}(\Omega)$.
The well-known fractional Poincaré inequality
$\|u\|_{L^{2}(\Omega)}\leq C[u]_{H^{s}({\mathbb{R}^{d}})},\qquad\forall u\in V,$ (1.4)
ensures that the seminorm $\cdot]_{H^{s}(\mathbb{R}^{d})}$ is in fact an equivalent norm on $V$. For further details, we refer the reader to [26, 13]. Finally, for $\alpha>0$ we denote by $W^{\alpha,\infty}(\Omega)$ the Hölder space of order $\alpha$
@@
Problem (1.1) can be written in weak form as follows:
@@
Remark 1.4.
Since we consider smooth domains $\Omega$ and piecewise linear basis functions defined over simplices, one must either over-mesh or under-mesh the domain. As the multiplication by $\delta^{s}$ disregards any information outside $\Omega$, it is technically more convenient to over-mesh.
@@
Let us also define $N_{h}\coloneq\dim V_{h}$. For the remaining elements, we drop the dependence on $h$ for the sake of the readability. We define $\{x_{i}\}_{i=1}^{N_{h}}$ be the vertices of all simplexes $K\in\mathcal{T}_{h}$ (including those lying in $\partial\Omega_{h}$). The canonical basis of $\mathrm{PL}(\mathcal{T}_{h})$ is given by the piece-wise linear function $\varphi_{i}$ such that $\varphi_{i}(x_{j})=\delta_{ij}$. The canonical basis of $V_{h}$ is $\phi_{i}=\delta^{s}\varphi_{i}$. When we use this basis, we will refer to the weighted Finite-Elements method, or WFEM for short.
@@
Let $s\in(0,1)$, $\partial\Omega\in C^{1,1}$ and $\delta$ satisfy ($\textup{A}_{\delta}^{\sigma}$) with $\sigma=1$. Then $V_{h}\subset V$.
@@
Due to Remark 1.3, if $\Omega=B_{1}$ and $f=1$, and we pick the distance $\delta(x)=(1-|x|^{2})_{+}$ we have precisely that $u/\delta^{s}$ is constant. This means that $u\in V_{h}$. Since $V_{h}\subset V$, then $u$ satisfies (1.12), and by uniqueness $u=u_{h}$. The numerical solution is exactly the true solution.
@@ 1.2 Main results
Let $s\in(0,1)$, $\partial\Omega\in C^{1,1}$, $\mu\in(s,2]$, and $\delta$ satisfy ($\textup{A}_{\delta}^{\sigma}$) with $\sigma=\max\{1,\mu\}$. Let also $u$ be the corresponding solution of (1.5), assume that $u/\delta^{s}\in W^{\mu,\infty}({\Omega})$, and $V_{h}$ be given by (1.11) and $u_{h}$ be the solution of (1.12). Then, we have that
@@
where $C$ does not depend on $h$. In particular, the maximal order $h^{2-s}\lvert\log h\rvert^{\frac{1}{2}}$ is achieved if $\mu=2$, i.e., $u/\delta^{s}\in W^{2,\infty}(\Omega)$.
@@
Let $s\in(0,1)$, $\gamma\in(s,3)$ such that $\gamma,\gamma\pm s\not\in\mathbb{N}$, $f\in W^{\gamma-s,\infty}(\Omega)$, $\partial\Omega\in C^{\gamma+1}\cap C^{1,1}$, and ($\textup{A}_{\delta}^{\sigma}$) holds with $\sigma=\gamma+1$. Let $u$ be the corresponding solution of (1.5), $V_{h}$ be given by (1.11), and $u_{h}$ be the solution of (1.12). Then, we have that
@@
Notice that in the results above, the constants $C$ depend on all the parameters in the statement $s,\mu$ (or $\gamma$), the geometric parameters $\Omega,d$, the constants in ($\textup{A}_{\delta}^{\sigma}$), and the uniform constants in the definition of the triangulation $\mathcal{T}_{h}$.
@@ 1.3 Comments on related literature
Our results indicate that the use of weighted basis can substantially improve the convergence rate, achieving up to $h^{2-s}\lvert\log h\rvert^{1/2}$, albeit at the cost of requiring higher regularity of both $f$ and $\Omega$. Furthermore, Theorem 4.9 in [5] demonstrates that when graded meshes are employed, the convergence rate in (1.15) can be improved up to an almost linear rate, proportional to $N^{-\frac{1}{d}}|\log N^{-\frac{1}{d}}|^{1/2}$, for $s\in(1/2,1)$ where in $N$ is the degrees of freedom (recall that in the case of uniform meshes $N_{h}\sim h^{-d}$). The study of $H^{1}$ rates is done in [11]. In the same paper the author show for dimension $d=1$ that graded meshes allow for $H^{s}$-rates of order $h^{2-s}$ (see [11, Remark 4.5]). In the present work, we piece-wise linearly approximate $u/\delta^{s}$ which sufficiently smooth everywhere in $\Omega$ so graded meshes do not seem to provide a major improvement. This also simplifies the implementation.
@@
We want to emphasize that both our work and the previously cited contributions address problem (1.1) involving the fractional Laplacian in its integral form, as defined in (1.2) over the entire space $\mathbb{R}^{d}$. However, the fractional Laplacian admits several distinct definitions (see, e.g., [43, 44]). Suitable localizations of these definitions to domains yield non-equivalent operators. Among these alternative formulations, the spectral fractional Laplacian has received particular attention due to its applications to problem (1.1) and to other equations involving fractional diffusion operators. The solutions of the corresponding Dirichlet problem behave like $u\sim d_{\Omega}$ ([17]), so the main difficulty of our framework is not present. This operator also admits a Caffarelli-Silvestre type extension ([55]) Finite element methods for the extension problem were first introduced in the pioneering work [46], achieving a convergence rate of $\lvert\log N\rvert^{s}N^{-1/(d+1)}$ in the $s$ fractional Sobolev norm corresponding to the problem, where $N$ stands for the degrees of freedom. To deal with computational challenges arising from the extra dimension, see [6]. Subsequently, in [7], the authors obtain exponential $H^{s}$ convergence rates using $hp$-FEM. For other approaches see [10, 24, 23, 56, 18].
We conclude by mentioning that specialised finite element bases have been employed in other contexts, such as acoustic and electromagnetic scattering, to mitigate the growth in the number of degrees of freedom required by standard finite element methods to maintain a given accuracy. We refer the reader to [19, 20], where products of piecewise polynomials with plane waves were chosen, or to [38], for oscillatory basis functions capturing the high-frequency solution behaviour in a transmission problem.
@@ 2 Proofs of the main results
Let $\Omega$ be such that $\partial\Omega\in C^{1,1}$. For any $\mu\in(0,2]$ consider a linear extension operator $E_{\mu}$ satisfying
@@
The existence of such an extension operator can be found in [54, Chapter VI, Theorem 3] and [36, Lemma 6.37]. We point out that in the construction of Stein $E_{\mu}=E_{1}$ for all $\mu\in(0,1]$. Furthermore, $E_{1}:C(\overline{\Omega})\to C_{b}({\mathbb{R}^{d}})$. Lastly, let us introduce the restriction $r_{\Omega}$ such that $r_{\Omega}u=u|_{\Omega}$. We first make the observation that for $C^{1,1}$ domains the power $s$ of the regularized distance function $\delta$ belongs to $V$.
@@
Assume that $\partial\Omega\in C^{1,1}$ and that $\delta\in W^{1,\infty}(\Omega)$ satisfies $c\delta\leq d_{\Omega}\leq C\delta$. Then, $\delta^{s}\in H^{s}({\mathbb{R}^{d}})$ for all $s\in(0,1)$.
@@
Since $\Omega\in C^{1,1}$ we can use the tubular neighbourhood theorem to prove that
$\int_{\Omega}d_{\Omega}(x)^{s-1}\mathrm{d}x<\infty.\qed$
@@ 2.1 Proof of convergence rates in $H^{s}$
As usual in the finite elements setting, we rely on the following Céa’s lemma to prove the desired convergence result.
@@
Let $s\in(0,1)$, $\partial\Omega\in C^{1,1}$ and $\delta$ satisfy ($\textup{A}_{\delta}^{\sigma}$) with $\sigma=1$. Let also $u$ be the corresponding solution of (1.5), $V_{h}$ be given by (1.11), and $u_{h}$ be the solution of (1.12). Then,
@@
We recall that $V_{h}\subset V$ as shown in Lemma 1.7. The $\geq$ inequality in (2.2) holds because $u_{h}\in V_{h}$. Let $v_{h}\in V_{h}$. For any $w_{h}\in V_{h}$ we can use the weak formulation to deduce $a(u-u_{h},w_{h})=a(u,w_{h})-a(u_{h},w_{h})=\int_{\Omega}fw_{h}-\int_{\Omega}fw_{h}=0$. Therefore, we have that
@@
In order to build a suitable competitor to apply Céa’s Lemma and obtain the desired orders of convergence, we need to define a suitable interpolant. More precisely, let $I_{h}:C(\mathbb{R}^{d})\to PL(\mathcal{T}_{h})\subset W^{1,\infty}({\Omega}_{h})$ be the piece-wise linear interpolation obtained by sampling on the vertexes of the triangles in $\mathcal{T}_{h}$. We define, for $\mu\in(s,2]$ the map $J_{h}:\delta^{s}W^{\mu,\infty}({\mathbb{R}^{d}})\to V_{h}$ by
@@
Applying the $K$-interpolation results in [45, Theorem 1.6, Example 1.8, result after the proof Example 1.8] that $T_{h}:W^{\alpha,\infty}({\mathbb{R}^{d}})\to W^{\alpha,\infty}({\mathbb{R}^{d}}),$ and we have
@@
First, we apply [22, Theorem 3.1.5] to show the result for some integer $\alpha,\mu$, i.e.,
@@
Let $s\in(0,1)$, $\partial\Omega\in C^{1,1}$, $\mu\in(s,2]$, and $\delta$ satisfy ($\textup{A}_{\delta}^{\sigma}$) with $\sigma=\max\{1,\mu\}$. Let also $v\in\delta^{s}C(\overline{\Omega})$ such that $v/\delta^{s}\in W^{\mu,\infty}(\overline{\Omega})$. Then, for all $\alpha\in(s,\min\{\mu,1\}]$, we have that
$v-J_{h}v]_{H^{s}(\mathbb{R}^{d})}\leq\frac{C_{1}}{\sqrt{\alpha-s}}h^{\mu-\alpha}\|v/\delta^{s}\|_{W^{\mu,\infty}({\Omega})},$ (2.7)
@@
$v-J_{h}v]_{H^{s}(\mathbb{R}^{d})}\leq C_{2}h^{\mu-s}\lvert\log h\rvert^{\frac{1}{2}}\|v/\delta^{s}\|_{W^{\mu,\infty}({\Omega})},$ (2.8)
@@
$v-J_{h}v=\delta^{s}E_{\alpha}r_{\Omega}\normalcolor E_{\mu}\frac{v}{\delta^{s}}-\delta^{s}E_{\alpha}r_{\Omega}\normalcolor I_{h}E_{\mu}\frac{v}{\delta^{s}}=\delta^{s}E_{\alpha}r_{\Omega}\normalcolor\left(E_{\mu}\frac{v}{\delta^{s}}-I_{h}E_{\mu}\frac{v}{\delta^{s}}\right)$